OpenStudy (anonymous):

Decompose -2x-23/2x^2-9x-5 into partial fractions

OpenStudy (anonymous):

\[\frac{ -2x-23 }{ 2x ^{2}-9x-5}\]

OpenStudy (anonymous):

i got \[\frac{ -3 }{ 2x+1 }+\frac{ 4 }{ x-5}\]

OpenStudy (anonymous):

How's it going

OpenStudy (anonymous):

??

OpenStudy (anonymous):

What I can't help you? I helped you before?

OpenStudy (anonymous):

yes, i gave my answer

OpenStudy (anonymous):

Its almost right

OpenStudy (anonymous):

When you use the system of equations, you find the coefficients of the partial fractions. A=-3 B=4 You switched your x-5 and 2x+1 equations and your A answer with -3

OpenStudy (anonymous):

So it would actually be \[-\frac{ 3 }{ x-5 } + \frac{ 4 }{ 2x+1 } instead of \frac{ -3}{ 2x+1 } + \frac{ 4}{ x-5 }\]

OpenStudy (anonymous):

i see where i went wrong, this really helped out a lot. thank you, sorry i didn't mean to seem rude earlier. i appreciate your help.

OpenStudy (anonymous):

No problem. Need any more help with any other questions you have?

OpenStudy (anonymous):

yes. im trying to solve this one \[\frac{ 14 }{ x ^{2}-3x }\]

OpenStudy (anonymous):

What kind method was provided

OpenStudy (anonymous):

wait it didnt go right..

OpenStudy (anonymous):

\[\frac{ 14 }{ x ^{2}-3x }-\frac{ 8 }{ x } >\frac{ -10 }{ x-3 }\]

OpenStudy (anonymous):

i got -19 <x<3 as a result

OpenStudy (anonymous):

Was there a certain method with it or is it decompose

OpenStudy (anonymous):

decompose

OpenStudy (anonymous):

You almost got that one right. You will be determining the given intervals to make each factor positive or negative. If the number of negative factors is odd, then the entire expression over this interval is negative. If the number of negative factors is even, then the entire expression over this interval is positive. x<-19 makes the expression NEGATIVE -19<x<0 makes the expression POSITIVE 0<x<3 makes the expression NEGATIVE 3<x makes the expression POSITIVE Since this is a greater than 0 inequality, all intervals that make the expression positive are part of the solution. -19 <x<0 or x>3

OpenStudy (anonymous):

Hows that?

OpenStudy (anonymous):

makes sense :3 im not too good with math and i usually make small mistakes. thank you

OpenStudy (anonymous):

Nice cat pictures btw.

OpenStudy (anonymous):

Medal for that lol :D